OpenStudy (austinl):

A 5.50 g bullet moving at 664 m/s strikes a 920 g wooden block at rest on a frictionless surface. The bullet emerges, traveling in the same direction with its speed reduced to 461 m/s. (a) What is the resulting speed of the block? (b) What is the speed of the bullet-block center of mass?

OpenStudy (carson889):

a) This is a conservation of momentum question. m_before * v_before = m_after * v_after. Let p1 be the momentum of the bullet Let p2 be the momentum of the block p1_initial + p2_initial = p1_final + p2_final \[p1_{i}=m_{i}*v_{i}=\frac{ 5.50g }{ 1000\frac{ g }{ kg } }*664 \frac{ m }{ s } = 3.65 kg\frac{ m }{ s }\] Final momentum of bullet: \[p1_{f}=m_{f}*v_{f}=\frac{ 5.50g }{ 1000\frac{ g }{ kg } }*461 \frac{ m }{ s } = 2.5355 kg\frac{ m }{ s }\] We know the initial momentum of the block was zero as its velocity was zero. Thus we have 3.65 + 0 = 2.5355 + m2*v2 1.1145 = m2*v2 We know m2 = 0.920 kg, so solve for v2 v2 = 1.211 m/s

OpenStudy (carson889):

v_com = (m1v1 + m2v2) / (m1 + m2) = (3.65 + 0) (kg m/s) / (.920 + 0.0055 ) kg = 3.944 m/s

OpenStudy (austinl):

Thank you very much!

OpenStudy (carson889):

You are welcome